# Presenting Interval Pomsets with Interfaces

### v1

@@ 4.1 Higher-dimensional automata
Languages of HDAs are closed under subsumption [10]: whenever $P\sqsubseteq Q\in L(X)$, also $P\in L(X)$.
@@ 4.2 ST-automata
An ST-automaton is a structure $A=(Q,E,I,F,\lambda)$ consisting of sets $Q$, $E\subseteq Q\times\Omega_{\simeq}\times Q$, $I,F\subseteq Q$, and a function $\lambda:Q\to\square$ such that for all $(q,{}_{S}U_{T},r)\in E$, $\lambda(q)=S$ and $\lambda(r)=T$.
@@ 4.3 From HDAs to ST-automata
For identities note that a path with a single cell $q$ is accepting in $X$ if and only if it is accepting in $F(X)$, and $\Phi(\textup{{id}}_{\textup{{ev}}(q)})=[\textup{{id}}_{\lambda(q)}]_{\sim}$. Now let $w=P_{1}\dots P_{m}\in L(F(X))$ be a non-identity. By definition, there exists $\pi=(q_{0},e_{1},q_{1},\dots,e_{n},q_{n})$ where $e_{i}=(q_{i-1},P^{\prime}_{i},q_{i})$, $P^{\prime}_{i}\in\Omega_{\simeq}$ such that $\textup{{id}}_{\lambda(q_{0})}P^{\prime}_{1}\textup{{id}}_{\lambda(q_{1})}\dots P^{\prime}_{n}\textup{{id}}_{\lambda(q_{n})}\sim w$. This means that $P^{\prime}_{1}*\dots*P^{\prime}_{n}$ is a decomposition of some $P\in L(X)$, hence $P\in\Phi(L(X))$. For the converse, let $w=P_{1}\dots P_{m}\in\Phi(L(X))$. Let $P^{\prime}_{1}*\dots*P^{\prime}_{n}$ be the sparse step decomposition of $P=P_{1}*\dots*P_{m}$. We have $P^{\prime}_{1}\dots P^{\prime}_{n}\sim w$. In addition, there exists an accepting path $\alpha=\beta_{1}*\dots*\beta_{n}$ in $X$ such that $\textup{{ev}}(\beta_{i})=P^{\prime}_{i}$. By construction there exists an accepting path $\pi=(\mathsf{src}(\beta_{1}),e_{1},\mathsf{tgt}(\beta_{1}),\dots,e_{n},\mathsf{tgt}(\beta_{n}))$ in $F(X)$ where $e_{i}=(\mathsf{src}(\beta_{i}),P^{\prime}_{i},\mathsf{tgt}(\beta_{i}))$. We have $\ell(\pi)\sim w$. ∎
@@ 4.4 From ST-automata to HDAs
{tikzpicture}
Figure 4: An ST-automaton with missing face relations, cf. Ex. 6.
@@
That is, each existing cell $q\in Q$ is associated with $(q,\emptyset,\emptyset)$ and $(q,B,C)$ is its (formal) face where $B$ is unstarted and $C$ terminated. Existing face maps $(q,r)\in\delta^{\nu}_{A,U}$ are copied to each pair $((q,B,C),(r,B,C))$ for which $A\subseteq U\setminus(B\cup C)$.
@@
Continuing Ex. 6, let again $A$ be the ST-automaton of Fig. 4. Then $G(A)$ is the HDA of Fig. 3 and $L(G(A))=\{bc\}\neq\emptyset=L(A)$.
@@
    for all $q\in Q$ with $\lambda(q)=\textup{{id}}_{U}$ and all $S,T\subseteq U$, there exist $p,r\in Q$ such that $(p,{}_{S}U_{U},q),(q,{}_{U}U_{T},r)\in E$;
@@ 5 Conclusion
Several previous works have studied interval pomsets with interfaces, their representations and their associated operational model. This paper unifies two different presentations (as a combinatorial object and as a word on a non-free monoid) and states how standard operations and transformations are expressed in both of these presentations.

### v2

@@ 4.1 Higher-dimensional automata
Languages of HDAs are closed under subsumption [11]: whenever $P\sqsubseteq Q\in L(X)$, then also $P\in L(X)$.
@@ 4.2 ST-automata
An ST-automaton is a structure $A=(Q,E,I,F,\lambda)$ consisting of sets $Q$, $E\subseteq Q\times\Omega_{\simeq}\times Q$, $I,F\subseteq Q$, and a function $\lambda:Q\to\square$ such that for all $(q,{\vphantom{U}}_{S}U_{T},r)\in E$, $\lambda(q)=S$ and $\lambda(r)=T$.
@@ 4.3 From HDAs to ST-automata
For identities note that a path with a single cell $q$ is accepting in $X$ if and only if it is accepting in $F(X)$, and $\Phi(\textup{{id}}_{\textup{{ev}}(q)})=[\textup{{id}}_{\lambda(q)}]_{\sim}$. Now let $w=P_{1}\dots P_{m}\in L(F(X))$ be a non-identity. By definition, there exists $\pi=(q_{0},e_{1},q_{1},\dots,e_{n},q_{n})$ where $e_{i}=(q_{i-1},P^{\prime}_{i},q_{i})$, $P^{\prime}_{i}\in\Omega_{\simeq}$ such that $\textup{{id}}_{\lambda(q_{0})}P^{\prime}_{1}\textup{{id}}_{\lambda(q_{1})}\dots P^{\prime}_{n}\textup{{id}}_{\lambda(q_{n})}\sim w$. This means that $P^{\prime}_{1}*\dots*P^{\prime}_{n}$ is a decomposition of some $P\in L(X)$, hence $w\in\Phi(L(X))$.
For the converse, let $w=P_{1}\dots P_{m}\in\Phi(L(X))$. Let $P^{\prime}_{1}*\dots*P^{\prime}_{n}$ be the sparse step decomposition of $P=P_{1}*\dots*P_{m}$. We have $P^{\prime}_{1}\dots P^{\prime}_{n}\sim w$. In addition, there exists an accepting path $\alpha=\beta_{1}*\dots*\beta_{n}$ in $X$ such that $\textup{{ev}}(\beta_{i})=P^{\prime}_{i}$. By construction there exists an accepting path $\pi=(\mathsf{src}(\beta_{1}),e_{1},\mathsf{tgt}(\beta_{1}),\dots,e_{n},\mathsf{tgt}(\beta_{n}))$ in $F(X)$ where $e_{i}=(\mathsf{src}(\beta_{i}),P^{\prime}_{i},\mathsf{tgt}(\beta_{i}))$. We have $\ell(\pi)\sim w$. ∎
@@ 4.4 From ST-automata to HDAs
[…]
Figure 4: An ST-automaton with missing face relations, cf. Ex. 7.
@@
That is, each existing cell $q\in Q$ is associated with $(q,\emptyset,\emptyset)$, and $(q,B,C)$ is its (formal) face where $B$ is unstarted and $C$ terminated. Existing face maps $(q,r)\in\delta^{\nu}_{A,U}$ are copied to each pair $((q,B,C),(r,B,C))$ for which $A\subseteq U\setminus(B\cup C)$.
@@
Continuing Ex. 7, let again $A$ be the ST-automaton of Fig. 4. Then $G(A)$ is the HDA of Fig. 3 and $L(G(A))=\{bc\}\neq\emptyset=L(A)$.
@@
    for all $q\in Q$ with $\lambda(q)=\textup{{id}}_{U}$ and all $S,T\subseteq U$, there exist $p,r\in Q$ such that $(p,{\vphantom{U}}_{S}U_{U},q),(q,{\vphantom{U}}_{U}U_{T},r)\in E$;
@@ 5 Conclusion
Several previous works have studied interval pomsets with interfaces, their representations, and their associated operational model. This paper unifies two different presentations (as a combinatorial object and as a word on a non-free monoid) and states how standard operations and transformations are expressed in both of these presentations.